OpenStudy (anonymous):

8 + 7y - y^2 answer: -(y+1)(y-8) how do you get the negative sign?

OpenStudy (anonymous):

In the original equation, they probably factored out a negative one before doing anything else, just to make the 7 positive.

OpenStudy (anonymous):

Is this correct? = -(y^2 + 7y + 8) = y^2 - 7y - 8 = y^2 + y - 8y - 8 = (y^2 + y) - (8y - 8)

OpenStudy (anonymous):

= y(y) - 8(y) =(y + 1)(y-8)

OpenStudy (anonymous):

I just dont know how they got the negative sign before the brackets....

OpenStudy (anonymous):

If you factor out a -, the 8 changes to -8

OpenStudy (anonymous):

?

OpenStudy (anonymous):

-8 times -1= 8 in your first one, its a +8, right? when you take out the -1, it turns into -8

OpenStudy (anonymous):

So that's how the negative sign was there....?

OpenStudy (anonymous):

You distribute

OpenStudy (anonymous):

e.g: